OpenStudy (anonymous):

What is the common ratio of the following sequence? 7, 21, 63, 189,…

OpenStudy (anonymous):

The common ratio means 189/63 = 63/21 = 21/7= 3 Do you understand? Cause this is pretty basic, if you dont understand this, it will kinda suck when you move on to tougher questions =(